OpenStudy (mitu12):

write -2.75 as a ratio of two integers

OpenStudy (anonymous):

\[-2.75 = - 2 \frac{ 3 }{ 4 } = - \frac{ 21 }{ 4 }\] and put the negative for any of them

OpenStudy (mitu12):

so there are both negative

OpenStudy (anonymous):

no.One only is negative \[\frac{ -21 }{ 4 } = \frac{ 21 }{ -4 }\]

OpenStudy (anonymous):

got me

OpenStudy (mitu12):

yes thank you

OpenStudy (anonymous):

it is 11 not 21 2 * 4 + 3 = 11 sorry!!

OpenStudy (mitu12):

ok thanks for telling me

OpenStudy (anonymous):

Hey do you know how to make \[a \frac{ b }{ c }= \frac{ e }{ c }\] e = a x c + b

OpenStudy (mitu12):

yes isnt it an improper fraction

OpenStudy (anonymous):

Yea right.I just wanted to make sure that you understood the solution.